2. Record the number of pins on the water. Be careful to not counts twice. There were __71__ pins in the water and ___14__ pins on the soil (non-water points). 3. Calculate a sample proportion of points that are on water and a 90% confidence interval for the proportion of the Earth that is covered with water. 90% Confidence Interval ________________________
Answer #1

n= 71 and x= 14

Standard error of proportion =

MARGIN OF ERROR= Zc*SE= 1.65*0.047= = 0.078

90% confidence interval is

90% confidence interval is  (0.119,0.275).